OpenStudy (redneckoutlaw):

Marks: 3 [10.01] Which is the simplified form of 16 times a to the eighth power all over 12 times a to the fourth power.? Choose one answer. a. 4 over 3. a4 b. 4a2 c. 4a4 d. 4 over 3. a2

jhonyy9 (jhonyy9):

do you think 16a^8 ------ = ? 12a^4 yes ?

jhonyy9 (jhonyy9):

separated 16 4*4 --- = ----- = ? 12 4*3 and a^8 ---- = a^(8-4) = ? a^4

jhonyy9 (jhonyy9):

so now just you need simplify

jhonyy9 (jhonyy9):

so the first fraction simplify by 4 and will remained 4/3 and down will get 8-4=4 so this mean a^4 than the right answer will be (4a^4)/3 hope so much that is understandably

jhonyy9 (jhonyy9):

what mean a. choise

jhonyy9 (jhonyy9):

ok ?